# Bidirectional Neural Networks for Global Nucleon-Nucleus Optical Model Calculations

### v1

@@ III Neural Network Architecture
For each combination of target and projectile type, I generate 100 random energy samples for each partial wave, yielding approximately 74,400 training examples in total. The wave functions are normalized by their maximum absolute value to bring all samples to a comparable scale. Each wave function is evaluated on the uniform grid of 100 radial points.
@@ IV Results
Figure 2: Comparison of wave functions predicted by BiLNN (dashed red lines) with numerically exact Numerov solutions (solid blue lines) for representative test cases. (a) p+208Pb at $E=83$ MeV, $l=0$: $s$-wave showing full penetration to the nuclear interior with $\sim$5 oscillations; (b) p+208Pb at $E=72$ MeV, $l=15$: high-$l$ partial wave with centrifugal barrier suppression at small $r$; (c) n+40Ca at $E=87$ MeV, $l=0$: neutron $s$-wave without Coulomb effects; (d) n+40Ca at $E=36$ MeV, $l=10$: lower-energy neutron with centrifugal suppression. The network accurately reproduces both amplitude and phase across all cases.
@@
The trained BiLNN is evaluated on a held-out test set comprising 20% of the total data (approximately 14,880 samples), with parameter combinations not seen during training. The overall relative error, defined as the root-mean-square deviation between predicted and true wave functions normalized by the wave function amplitude, is 1.2% averaged over all test samples. This level of accuracy is sufficient for most practical applications, as it corresponds to errors in the scattering $S$-matrix of comparable magnitude.
@@
Figure 2 presents a detailed comparison of wave functions predicted by BiLNN with numerically exact Numerov solutions for four representative test cases that span different projectile types, target nuclei, energies, and partial waves. Panel (a) shows the $s$-wave ($l=0$) radial wave function for proton scattering on 208Pb at $E=83$ MeV, where the real part exhibits approximately 5 complete oscillations within the radial range r=0r=0–15 fm, with amplitude growing from small values near the origin (where the strong nuclear and Coulomb potentials suppress the wave function) to maximum amplitude around $r\approx 8$ fm before settling into the asymptotic Coulomb behavior. The BiLNN prediction (dashed red) overlays the Numerov reference (solid blue) with deviations barely visible on this scale, demonstrating sub-percent accuracy in both amplitude and phase. Panel (b) displays the $l=15$ partial wave for the same p+208Pb system at $E=72$ MeV. Here the centrifugal barrier $l(l+1)/r^{2}$ creates a classically forbidden region at small radii, suppressing the wave function to nearly zero for $r\lesssim 5$ fm. Beyond the classical turning point, the wave function emerges with oscillatory behavior confined to the exterior region $r>5$ fm. The network accurately captures both the exponential suppression in the forbidden region and the subsequent oscillatory structure, including the correct phase at the matching radius. Panels (c) and (d) show neutron scattering on 40Ca, where the absence of Coulomb interaction simplifies the potential but the wave function must still satisfy the correct asymptotic matching to spherical Bessel functions. At $E=87$ MeV [panel (c), $l=0$], the $s$-wave penetrates to the origin and shows approximately 5 oscillations with the characteristic pattern of increasing amplitude toward the asymptotic region. At $E=36$ MeV [panel (d), $l=10$], the lower energy combined with the centrifugal barrier produces a wave function that is suppressed for $r\lesssim 4$ fm and exhibits only a single oscillation in the exterior region, with longer wavelength reflecting the lower kinetic energy. In all cases, the phase agreement between BiLNN and Numerov is excellent across the entire radial range, demonstrating that the phase-space coordinate representation successfully enables generalization across the wide energy range.
@@
Figure 3: $S$-matrix comparison for p+208Pb scattering at three energies: $E=27$ MeV (left), $E=83$ MeV (center), and $E=157$ MeV (right). Top row [(a)–(c)]: $|S_{l}|$ showing oscillatory absorption at low $l$ with sharp transition to transparency ($|S_{l}|\to 1$) at the grazing angular momentum $l_{g}$, which increases with energy from $l_{g}\approx 12$ at 27 MeV to $l_{g}\approx 20$ at 157 MeV. Bottom row [(d)–(f)]: $\arg(S_{l})$ showing the nuclear and Coulomb phase shifts, with rapid $l$-dependence at low energies that smooths out at higher energies. Blue circles: Numerov reference; red squares: BiLNN prediction.
@@
Since the ultimate goal of optical model calculations is to predict scattering observables, I verify that the wave function accuracy translates to accurate $S$-matrix elements and cross sections. The $S$-matrix is extracted from the predicted wave functions using the amplitude integral method [4]
$f_{l}=-\frac{1}{E}\int_{0}^{\infty}F_{l}(\eta,kr)\,V(r)\,\psi_{l}(r)\,dr,$ (12)
where $F_{l}$ is the regular Coulomb function, giving $S_{l}=1+2ikf_{l}$. This integral formulation is preferred over asymptotic matching because it weights the wave function by the potential, making it less sensitive to errors in the asymptotic region where the wave function amplitude is large but the physical information content is low.
@@
Figure 3 compares the $S$-matrix modulus $|S_{l}|$ and phase $\arg(S_{l})$ obtained from BiLNN and Numerov wave functions for proton scattering on 208Pb at three representative energies: 27 MeV (left column), 83 MeV (center column), and 157 MeV (right column). The top row [panels (a)–(c)] displays $|S_{l}|$ as a function of partial wave $l$. At the lowest energy $E=27$ MeV [panel (a)], the $S$-matrix modulus shows a complex oscillatory pattern at low $l$, with $|S_{l}|$ varying between 0.4 and 0.75 for $l\lesssim 10$, reflecting interference between nuclear and Coulomb amplitudes in this strongly absorbing regime. Beyond $l\approx 12$, the transition to transparency ($|S_{l}|\approx 1$) is sharp, occurring over just 2–3 partial waves. At intermediate energy $E=83$ MeV [panel (b)], the absorption region extends to higher partial waves ($l\lesssim 14$), with $|S_{l}|$ showing similar oscillatory structure in the interior region before the sharp rise to unity. At the highest energy $E=157$ MeV [panel (c)], the grazing angular momentum shifts further to $l_{g}\approx 20$, and the oscillatory pattern in $|S_{l}|$ becomes more pronounced, with values dropping as low as 0.3 before recovering. This energy-dependent shift of the absorption edge reflects the increasing classical impact parameter $b=l/k$ at higher energies, with the grazing condition $b\approx R$ determining where the projectile trajectory begins to miss the nuclear volume. The BiLNN predictions (red squares) overlay the Numerov reference values (blue circles) with excellent agreement across all partial waves, accurately capturing the oscillatory structure, absorption depth, and transition sharpness. The bottom row [panels (d)–(f)] shows the $S$-matrix phase $\arg(S_{l})$, which encodes the combined nuclear and Coulomb phase shifts. At low energies [panel (d)], the phase exhibits rapid variation with $l$, starting near +2+2 radians at $l=0$ and dropping sharply to 2-2 radians around $l\approx 6$ before recovering to zero at high $l$. At higher energies [panels (e) and (f)], the phase variation becomes smoother but extends to higher $l$ values before asymptoting to zero. The network reproduces these phase patterns with high fidelity, which is crucial since even small phase errors would produce significant deviations in the interference patterns of the cross section.
@@
Figure 4: Proton elastic scattering angular distributions $d\sigma/d\sigma_{\text{Ruth}}$ for three target nuclei at two energies each. (a,d) 40Ca at $E=27$ MeV and $E=97$ MeV; (b,e) 90Zr at $E=31$ MeV and $E=56$ MeV; (c,f) 208Pb at $E=27$ MeV and $E=83$ MeV. The top row shows lower-energy cases with broader diffraction features; the bottom row shows higher-energy cases with more compressed, sharper minima. Blue solid lines: Numerov reference; red dashed lines: BiLNN prediction. The diffraction patterns, including positions, depths, and widths of multiple minima, are reproduced across up to four orders of magnitude in cross section ratio.
@@
Figure 5: Neutron elastic scattering angular distributions $d\sigma/d\Omega$ for six target nuclei spanning the training set. (a) 12C at $E=49$ MeV; (b) 40Ca at $E=34$ MeV; (c) 90Zr at $E=22$ MeV; (d) 56Fe at $E=83$ MeV; (e) 120Sn at $E=74$ MeV; (f) 208Pb at $E=63$ MeV. Blue solid lines: Numerov reference; red dashed lines: BiLNN prediction.
@@
Figure 6 provides a comprehensive summary of the error distribution across the parameter space, restricted to physically contributing partial waves satisfying $l<kR+5$ where $R\approx 1.2A^{1/3}$ fm. This criterion excludes high-$l$ partial waves for light nuclei at low energies, where the centrifugal barrier prevents the wave function from penetrating into the nuclear interior and therefore makes negligible contribution to the cross section. Panel (a) displays the relative error as a function of partial wave quantum number $l$. The error shows a non-monotonic dependence on angular momentum: both low-$l$ partial waves ($l\lesssim 4$) and high-$l$ partial waves ($l\gtrsim 24$) exhibit errors of approximately 1.5–1.8%, while intermediate partial waves ($l\approx 8$–16) achieve the lowest errors around 1.0%. The elevated error at low $l$ can be understood physically: for $s$-waves and low partial waves, the wave function penetrates deep into the nuclear interior where both the nuclear attraction and Coulomb repulsion (for protons) are strongest, requiring the network to capture rapid variations in a region of steep potential gradients. The slightly elevated error at high $l$ may reflect the increased sensitivity to the asymptotic matching region, where small phase errors accumulate over many oscillations. Panel (b) shows the error breakdown by target nucleus for all twelve nuclei in the training set. The errors are uniform for light and medium-mass nuclei from 12C to 120Sn, clustering around 0.8–1.3%, while the heaviest targets 197Au and 208Pb show somewhat larger errors of 1.6–1.8%. This trend can be attributed to the stronger Coulomb interaction for proton scattering on heavy targets, which introduces additional oscillatory structure at large distances and requires more extended matching radii. The horizontal dashed line indicates the overall average error of 1.2%. Panel (c) reveals the most pronounced systematic dependence: the error varies significantly with projectile energy, from approximately 4% at the lowest energies (1–20 MeV) down to about 1% at high energies (above 100 MeV). This trend reflects the challenge of learning wave functions at low energies where the de Broglie wavelength is longest: the wave function exhibits fewer oscillations within the nuclear interior, providing less structure for the network to learn, and the relative phase error has larger impact on the overall solution. At higher energies, the more rapid oscillations are well captured by the phase-space coordinate representation, resulting in errors close to or below the overall average.
@@
Figure 6: Summary of neural network prediction errors for physically contributing partial waves ($l<kR+5$, where $R\approx 1.2A^{1/3}$ fm). (a) Error versus partial wave $l$, showing elevated errors ($\sim$1.5–1.8%) at both low $l$ (deep nuclear penetration) and high $l$ (asymptotic sensitivity), with minimum errors ($\sim$1%) at intermediate $l\approx 8$–16; (b) Error by target nucleus for all twelve training targets, with uniform accuracy (0.8–1.3%) for light-to-medium nuclei and slightly higher errors (1.6–1.8%) for the heaviest targets; (c) Error versus projectile energy, showing the strongest systematic dependence: $\sim$4% at low energies (1–20 MeV) decreasing to $\sim$1% above 100 MeV. The dashed lines indicate the overall average error of 1.2%.
@@
A key test of whether the network has learned the underlying physics rather than merely memorizing the training data is its ability to extrapolate to nuclei not included in training. Figure 7 shows predictions for three nuclei that were never seen during training: 24Mg ($A=24$, $Z=12$), 63Cu ($A=63$, $Z=29$), and 184W ($A=184$, $Z=74$). These targets were deliberately chosen to lie between the training nuclei in atomic mass: 24Mg between 16O and 27Al, 63Cu between 58Ni and 90Zr, and 184W between 120Sn and 197Au, testing interpolation capability with nuclei having different proton-to-neutron ratios that affect the optical potential parameters through the isospin dependence of the KD02 parameterization. The top row [panels (a)–(c)] displays $s$-wave ($l=0$) proton scattering wave functions at intermediate energies. For 24Mg at $E=76$ MeV [panel (a)], the light target produces a wave function with approximately 4 oscillations and relatively uniform amplitude, characteristic of weak absorption in the small nuclear volume. For 63Cu at $E=88$ MeV [panel (b)], the larger nuclear radius shifts the oscillatory pattern outward with approximately 5 oscillations visible, while the wave function amplitude grows more pronounced toward the asymptotic region. For 184W at $E=38$ MeV [panel (c)], the combination of heavy target and lower energy produces a wave function with fewer oscillations (longer wavelength) but still excellent agreement between BiLNN and Numerov across the full radial range. The bottom row [panels (d)–(f)] shows the corresponding elastic scattering cross sections $d\sigma/d\sigma_{\text{Ruth}}$. For 24Mg [panel (d)], the cross section ratio remains above unity across most angles with a single broad minimum. For 63Cu [panel (e)], multiple diffraction minima appear spanning nearly four orders of magnitude in cross section ratio. For 184W [panel (f)], the strong Coulomb field produces a pattern with the ratio declining smoothly from forward angles to backward angles. In all cases, the BiLNN predictions closely track the Numerov reference, demonstrating that the network has learned the smooth dependence of the scattering solution on atomic mass $A$ and charge $Z$ embedded in the optical potential, rather than memorizing target-specific patterns.
@@
Figure 7: Extrapolation test for nuclei not included in training: 24Mg (left), 63Cu (center), and 184W (right). Top row [(a)–(c)]: $s$-wave proton wave functions at $E=76$, 88, and 38 MeV, respectively, showing the real part of $\psi_{l}(r)$; bottom row [(d)–(f)]: corresponding elastic scattering cross sections $d\sigma/d\sigma_{\text{Ruth}}$ spanning up to four orders of magnitude. Blue: Numerov reference; red: BiLNN prediction. The network generalizes successfully to these unseen nuclei with accuracy comparable to the training set, demonstrating that it has learned the physics of the optical model rather than memorizing target-specific patterns.

### v2

@@ III Neural Network Architecture
For each combination of target and projectile type, I generate 200 random energy samples, with all partial waves computed at each energy, yielding approximately 148,800 training examples in total. The wave functions are normalized by their maximum absolute value to bring all samples to a comparable scale. Each wave function is evaluated on the uniform grid of 100 radial points.
@@ IV Results
Figure 2: Comparison of wave functions predicted by BiLNN (dashed red lines) with numerically exact Numerov solutions (solid blue lines) for representative test cases. (a) p+208Pb at $E=75$ MeV, $l=0$: $s$-wave showing full penetration to the nuclear interior with $\sim$5 oscillations; (b) p+208Pb at $E=75$ MeV, $l=15$: high-$l$ partial wave with centrifugal barrier suppression at small $r$; (c) n+40Ca at $E=69$ MeV, $l=0$: neutron $s$-wave without Coulomb effects; (d) n+40Ca at $E=69$ MeV, $l=10$: centrifugal suppression at intermediate energy. The network accurately reproduces both amplitude and phase across all cases.
@@
The trained BiLNN is evaluated on a held-out test set comprising 20% of the total data (approximately 29,760 samples), with parameter combinations not seen during training. The overall relative error, defined as the root-mean-square deviation between predicted and true wave functions normalized by the wave function amplitude, is 1.2% averaged over all test samples. The corresponding $S$-matrix error, computed using the matching method averaged over 15 radii, is approximately 0.6%, lower than the wave function error because $S$-matrix extraction is less sensitive to local wave function details.
@@
Figure 2 presents a detailed comparison of wave functions predicted by BiLNN with numerically exact Numerov solutions for four representative test cases that span different projectile types, target nuclei, energies, and partial waves. Panel (a) shows the $s$-wave ($l=0$) radial wave function for proton scattering on 208Pb at $E=75$ MeV, where the real part exhibits approximately 5 complete oscillations within the radial range r=0r=0–15 fm, with amplitude growing from small values near the origin (where the strong nuclear and Coulomb potentials suppress the wave function) to maximum amplitude around $r\approx 8$ fm before settling into the asymptotic Coulomb behavior. The BiLNN prediction (dashed red) overlays the Numerov reference (solid blue) with deviations barely visible on this scale, demonstrating sub-percent accuracy in both amplitude and phase. Panel (b) displays the $l=15$ partial wave for the same p+208Pb system at $E=75$ MeV. Here the centrifugal barrier $l(l+1)/r^{2}$ creates a classically forbidden region at small radii, suppressing the wave function to nearly zero for $r\lesssim 5$ fm. Beyond the classical turning point, the wave function emerges with oscillatory behavior confined to the exterior region $r>5$ fm. The network accurately captures both the exponential suppression in the forbidden region and the subsequent oscillatory structure, including the correct phase at the matching radius. Panels (c) and (d) show neutron scattering on 40Ca, where the absence of Coulomb interaction simplifies the potential but the wave function must still satisfy the correct asymptotic matching to spherical Bessel functions. At $E=69$ MeV [panel (c), $l=0$], the $s$-wave penetrates to the origin and shows approximately 4 oscillations with the characteristic pattern of increasing amplitude toward the asymptotic region. At $E=69$ MeV [panel (d), $l=10$], the centrifugal barrier produces a wave function that is suppressed for $r\lesssim 4$ fm and exhibits oscillatory behavior in the exterior region. In all cases, the phase agreement between BiLNN and Numerov is excellent across the entire radial range, demonstrating that the phase-space coordinate representation successfully enables generalization across the wide energy range.
@@
Figure 3: $S$-matrix comparison for p+208Pb scattering at three energies: $E=18$ MeV (left), $E=75$ MeV (center), and $E=197$ MeV (right). Top row [(a)–(c)]: $|S_{l}|$ showing oscillatory absorption at low $l$ with sharp transition to transparency ($|S_{l}|\to 1$) at the grazing angular momentum $l_{g}$, which increases with energy from $l_{g}\approx 8$ at 18 MeV to $l_{g}\approx 22$ at 197 MeV. Bottom row [(d)–(f)]: $\arg(S_{l})$ showing the nuclear and Coulomb phase shifts, with rapid $l$-dependence at low energies that smooths out at higher energies. Blue circles: Numerov reference; red squares: BiLNN prediction.
@@
Since the ultimate goal of optical model calculations is to predict scattering observables, I verify that the wave function accuracy translates to accurate $S$-matrix elements and cross sections. The $S$-matrix is extracted from the predicted wave functions using the asymptotic matching method [20]
$S_{l}=\frac{H_{l}^{(-)^{\prime}}\psi_{l}-H_{l}^{(-)}\psi_{l}^{\prime}}{H_{l}^{(+)^{\prime}}\psi_{l}-H_{l}^{(+)}\psi_{l}^{\prime}},$ (12)
where $H_{l}^{(\pm)}$ are the Coulomb-Hankel functions and $\psi_{l}$, $\psi_{l}^{\prime}$ are the wave function and its derivative evaluated at the matching radius. This ratio formulation is preferred because it is independent of the wave function normalization: any arbitrary scaling factor cancels between numerator and denominator. This property is essential because the neural network is trained on wave functions normalized to $\max|\psi_{l}(r)|=1$ for numerical stability, rather than the physical normalization $\psi_{l}\to(H_{l}^{(-)}-S_{l}H_{l}^{(+)})/(2ik)$ as $r\to\infty$. The matching method extracts the correct $S$-matrix regardless of this arbitrary normalization choice.
To reduce numerical noise from the derivative estimation at a single radius, I average the extracted $S$-matrix over 15 consecutive matching radii in the asymptotic region (r=12r=12–15 fm), where the nuclear potential has decayed to negligible values. This multi-point averaging significantly improves numerical stability without introducing systematic bias, since all radii in this region should yield identical $S$-matrix values for the exact solution.
The error propagation from wave function to observables exhibits an important hierarchy. The 1.2% wave function error translates to only 0.6% $S$-matrix error because the matching method involves ratios that partially cancel local errors. However, the differential cross section $d\sigma/d\Omega$ involves a coherent sum over $\sim$30 partial waves,
$\frac{d\sigma}{d\Omega}=\left|f_{C}(\theta)+\sum_{l}\frac{2l+1}{2ik}(S_{l}-1)P_{l}(\cos\theta)e^{2i\sigma_{l}}\right|^{2},$ (13)
where small phase errors in individual $S_{l}$ can accumulate. At forward angles where constructive interference dominates, cross-section errors remain modest ($\sim$1–10%). At backward angles ($\theta\approx 180\textdegree$), where the Coulomb amplitude $f_{C}$ and nuclear amplitude $f_{N}$ have comparable magnitudes but opposite phases, destructive interference leads to catastrophic cancellation: the total amplitude $|f_{C}+f_{N}|\ll|f_{C}|,|f_{N}|$ becomes highly sensitive to small phase errors, amplifying the cross-section error to $\sim$50%. This is a fundamental numerical sensitivity of backward-angle scattering, not a limitation specific to the neural network. Integral quantities such as the reaction cross section $\sigma_{R}=\pi\lambda^{2}\sum_{l}(2l+1)(1-|S_{l}|^{2})$, which do not involve coherent interference, maintain errors below 0.1%.
@@
Figure 3 compares the $S$-matrix modulus $|S_{l}|$ and phase $\arg(S_{l})$ obtained from BiLNN and Numerov wave functions for proton scattering on 208Pb at three representative energies: 18 MeV (left column), 75 MeV (center column), and 197 MeV (right column). The top row [panels (a)–(c)] displays $|S_{l}|$ as a function of partial wave $l$. At the lowest energy $E=18$ MeV [panel (a)], the $S$-matrix modulus shows a complex oscillatory pattern at low $l$, with $|S_{l}|$ varying between 0.4 and 0.9 for $l\lesssim 8$, reflecting interference between nuclear and Coulomb amplitudes in this strongly absorbing regime. Beyond $l\approx 8$, the transition to transparency ($|S_{l}|\approx 1$) is sharp, occurring over just 2–3 partial waves. At intermediate energy $E=75$ MeV [panel (b)], the absorption region extends to higher partial waves ($l\lesssim 14$), with $|S_{l}|$ showing oscillatory structure in the interior region before the sharp rise to unity. At the highest energy $E=197$ MeV [panel (c)], the grazing angular momentum shifts further to $l_{g}\approx 22$, and the oscillatory pattern in $|S_{l}|$ becomes more pronounced, with values dropping as low as 0.2 before recovering. This energy-dependent shift of the absorption edge reflects the increasing classical impact parameter $b=l/k$ at higher energies, with the grazing condition $b\approx R$ determining where the projectile trajectory begins to miss the nuclear volume. The BiLNN predictions (red squares) overlay the Numerov reference values (blue circles) with excellent agreement across all partial waves, accurately capturing the oscillatory structure, absorption depth, and transition sharpness. The bottom row [panels (d)–(f)] shows the $S$-matrix phase $\arg(S_{l})$, which encodes the combined nuclear and Coulomb phase shifts. At low energies [panel (d)], the phase exhibits rapid variation with $l$, starting near +2+2 radians at $l=0$ and dropping sharply to 2-2 radians around $l\approx 6$ before recovering to zero at high $l$. At higher energies [panels (e) and (f)], the phase variation becomes smoother but extends to higher $l$ values before asymptoting to zero. The network reproduces these phase patterns with high fidelity, which is crucial since even small phase errors would produce significant deviations in the interference patterns of the cross section.
@@
Figure 4: Proton elastic scattering angular distributions $d\sigma/d\sigma_{\text{Ruth}}$ for three target nuclei at two energies each. (a,d) 40Ca at $E=30$ MeV and $E=65$ MeV; (b,e) 90Zr at $E=30$ MeV and $E=65$ MeV; (c,f) 208Pb at $E=30$ MeV and $E=65$ MeV. The top row shows lower-energy cases with broader diffraction features; the bottom row shows higher-energy cases with more compressed, sharper minima. Blue solid lines: Numerov reference; red dashed lines: BiLNN prediction. The diffraction patterns, including positions, depths, and widths of multiple minima, are reproduced across up to four orders of magnitude in cross section ratio.
@@
Figure 5: Neutron elastic scattering angular distributions $d\sigma/d\Omega$ for six target nuclei spanning the training set. (a) 12C at $E=30$ MeV; (b) 40Ca at $E=30$ MeV; (c) 90Zr at $E=30$ MeV; (d) 56Fe at $E=65$ MeV; (e) 120Sn at $E=65$ MeV; (f) 208Pb at $E=65$ MeV. Blue solid lines: Numerov reference; red dashed lines: BiLNN prediction.
@@
Figure 6 provides a comprehensive summary of the $S$-matrix error distribution across the parameter space, restricted to physically contributing partial waves satisfying $l<kR+5$ where $R\approx 1.2A^{1/3}$ fm. This criterion excludes high-$l$ partial waves for light nuclei at low energies, where the centrifugal barrier prevents the wave function from penetrating into the nuclear interior and therefore makes negligible contribution to the cross section. The $S$-matrix error is computed as the relative error in $|S_{l}|$ between BiLNN and Numerov, averaged over 15 matching radii to ensure numerical stability. Panel (a) displays the relative error as a function of partial wave quantum number $l$. The error shows a non-monotonic dependence on angular momentum: low-$l$ partial waves ($l\lesssim 4$) exhibit the largest errors ($\sim$1.0–1.3%), while higher partial waves achieve progressively lower errors ($\lesssim$0.5%). The elevated error at low $l$ can be understood physically: for $s$-waves and low partial waves, the wave function penetrates deep into the nuclear interior where both the nuclear attraction and Coulomb repulsion (for protons) are strongest, requiring the network to capture rapid variations in a region of steep potential gradients. Panel (b) shows the error breakdown by target nucleus for all twelve nuclei in the training set. The errors are uniform across most nuclei, clustering around 0.3–0.8%, with the heaviest targets 197Au and 208Pb showing slightly larger errors around 1%. This trend can be attributed to the stronger Coulomb interaction for proton scattering on heavy targets, which introduces additional oscillatory structure at large distances and requires more extended matching radii. The horizontal dashed line indicates the overall average $S$-matrix error of 0.6%. Panel (c) reveals the most pronounced systematic dependence: the error varies significantly with projectile energy, from approximately 2.5% at the lowest energies (10–20 MeV) down to about 0.2–0.5% at high energies (above 50 MeV). This trend reflects the challenge of learning wave functions at low energies where the de Broglie wavelength is longest: the wave function exhibits fewer oscillations within the nuclear interior, providing less structure for the network to learn, and the relative phase error has larger impact on the overall solution. At higher energies, the more rapid oscillations are well captured by the phase-space coordinate representation, resulting in $S$-matrix errors well below 1%.
@@
Figure 6: Summary of $S$-matrix prediction errors for physically contributing partial waves ($l<kR+5$, where $R\approx 1.2A^{1/3}$ fm). (a) Error versus partial wave $l$, showing elevated errors ($\sim$1.0–1.3%) at low $l$ (deep nuclear penetration), decreasing to $\lesssim$0.5% at higher $l$; (b) Error by target nucleus for all twelve training targets, with uniform accuracy (0.3–0.8%) across most nuclei and slightly higher errors ($\sim$1%) for the heaviest targets; (c) Error versus projectile energy, showing the strongest systematic dependence: $\sim$2.5% at low energies (10–20 MeV) decreasing to $\lesssim$0.5% above 50 MeV. The dashed lines indicate the overall average $S$-matrix error of 0.6%.
@@
A key test of whether the network has learned the underlying physics rather than merely memorizing the training data is its ability to extrapolate to nuclei not included in training. Figure 7 shows predictions for three nuclei that were never seen during training: 24Mg ($A=24$, $Z=12$), 63Cu ($A=63$, $Z=29$), and 184W ($A=184$, $Z=74$). These targets were deliberately chosen to lie between the training nuclei in atomic mass: 24Mg between 16O and 27Al, 63Cu between 58Ni and 90Zr, and 184W between 120Sn and 197Au, testing interpolation capability with nuclei having different proton-to-neutron ratios that affect the optical potential parameters through the isospin dependence of the KD02 parameterization. The top row [panels (a)–(c)] displays $s$-wave ($l=0$) proton scattering wave functions at $E=50$ MeV. For 24Mg [panel (a)], the light target produces a wave function with approximately 4 oscillations and relatively uniform amplitude, characteristic of weak absorption in the small nuclear volume. For 63Cu [panel (b)], the larger nuclear radius shifts the oscillatory pattern outward with approximately 5 oscillations visible, while the wave function amplitude grows more pronounced toward the asymptotic region. For 184W [panel (c)], the heavy target produces a wave function with similar oscillation count but distinct amplitude pattern due to stronger absorption, with excellent agreement between BiLNN and Numerov across the full radial range. The bottom row [panels (d)–(f)] shows the corresponding elastic scattering cross sections $d\sigma/d\sigma_{\text{Ruth}}$. For 24Mg [panel (d)], the cross section ratio remains above unity across most angles with a single broad minimum. For 63Cu [panel (e)], multiple diffraction minima appear spanning nearly four orders of magnitude in cross section ratio. For 184W [panel (f)], the strong Coulomb field produces a pattern with the ratio declining smoothly from forward angles to backward angles. In all cases, the BiLNN predictions closely track the Numerov reference, demonstrating that the network has learned the smooth dependence of the scattering solution on atomic mass $A$ and charge $Z$ embedded in the optical potential, rather than memorizing target-specific patterns.
@@
Figure 7: Extrapolation test for nuclei not included in training: 24Mg (left), 63Cu (center), and 184W (right). Top row [(a)–(c)]: $s$-wave proton wave functions at $E=50$ MeV, showing the real part of $\psi_{l}(r)$; bottom row [(d)–(f)]: corresponding elastic scattering cross sections $d\sigma/d\sigma_{\text{Ruth}}$ spanning up to four orders of magnitude. Blue: Numerov reference; red: BiLNN prediction. The network generalizes successfully to these unseen nuclei with accuracy comparable to the training set, demonstrating that it has learned the physics of the optical model rather than memorizing target-specific patterns.